The next dividend for GTA2 corp will be $4 per share. Investors require a 16% return. Dividends of GTA2 increases by 6% every year. Based on this information what is the value of the stock today? (constant growth model) , what is the value of the stock 4 years from today ? 100 years from today?
Answer #1

Value of Stock is equal to the present value of all future dividends

= Expected Dividend/(Required Return - growth rate)

= 4/(16%-6%)

= $40

Value of stock after 4 years = Dividend in year 5/(Required Return – Growth Rate)

= 4(1.06)4/10%

= $50.50

100 years from today = 4(1.06)100/10%

= 13,572.08